You and your friend are applying to speak at graduation. If there are 16 total entries to speak at graduation and two people are chosen at random, what is the probability that you and your friend are chosen?


1
Expert's answer
2021-02-18T12:38:32-0500

Total entries are 16 and number of persons chosen at random is two. Thus the total events can be possible are "16*15=240"

i.e. 16 type of persons can be come on first randomly choose and due to non repetition of

persons there are only 15 different people can be choose on second time.


The favourable events are 2

i.e. me on first call and my friend on second call and my friend is on first call and me on second call.

"p"(me and my friend) "=" ( favourable event)/(total event) "=" "\frac{2}{240}=0.008333"